# Quantum-enhanced mean value estimation via adaptive measurement

### v3

@@ 3.4 The proposed adaptive estimation algorithm
Our algorithm is based on the maximum likelihood estimation method for the random variables subjected to the probability distribution (3.3). An overview is shown in Fig. 1. The algorithm is composed of the following four steps.
@@
    Identify the probabilities of depolarization $1-p_{\rm q}$ by some experiments on the quantum device to be used. For the first measurement, we set the amplified level as $\alpha_{1}=1$.
The impact of calibration errors of $p_{\rm q}$ on this algorithm is analyzed numerically in Appendix F. In the following, we assume the calibration is perfect unless otherwise specified. Unlike Bayesian estimation, our algorithm does not require any prior information on the target value.
@@
    Measure the quantum state (15) $N$ times independently with either POVM (3.3) or (3.3), according to the amplified level $\alpha_{k}$. We write the total number of hitting "1" as $x^{(k)}\in\{0,1,\cdots,N\}$. Here, $x^{(k)}$ is a realization of the Binomial random variable
@@
Note that the measurement $\bm{x}_{k}$ is the collection of all the results via the measurements with $\alpha_{l}~{}\forall l=1,2,\cdots,k$. Since the amplified level $\alpha_{k}$ depends on the previous results $\bm{x}_{k-1}$ as described below (except for the case $k=1$), the likelihood function $\mathcal{L}_{k}(\theta;\bm{x}_{k})$ has a hierarchical structure, unlike the case of Eq. (6).
@@
    To efficiently read out $\theta^{*}$ in the sense of Fisher information, determine $\alpha_{k+1}$ for the next measurement based on the current estimate $\hat{\theta}_{k}$ by numerically solving the following optimization problem:
@@
    where $D_{k}\subset\mathbb{N}$ is a (finite) optimization range and $\delta\in(0,1]$ is a parameter for regularization. Then return to (ii) and perform the measurement again.
The optimization problem (35) comes from the third requirement described in the beginning of this subsection; we will explain a detailed explanation in the next two paragraphs. Since the elements of $D_{k}$ are integers, the optimization problem can be efficiently solved using a simple brute force method. Also, it is obvious that the optimization does not yield an amplified level such that $\sin{(\alpha\hat{\theta}_{k})}\simeq 0$, and thus our method can avoid vanishing classical Fisher information regardless of the target value.
@@
Although the optimization seems to be desirable, we actually encounter the following issue. That is, if $\kappa_{\infty}$ is too close to one, the probability $1-\mathbf{P}_{\mathcal{D}}(\alpha;\theta^{*})$ becomes small for such an $\alpha$ when $d\gg 1$, meaning that we need a large number of shots $N$ to have a good estimate. To circumvent this issue, we introduce a regularization function ${\rm Reg}(\alpha;\hat{\theta}_{k})$ such that ${\rm Reg}\simeq 0$ at $\kappa_{\infty}\simeq 1$ and take the objective function $\kappa_{\infty}\mathcal{I}_{\rm q}/\alpha\times{\rm Reg}$ rather than $\kappa_{\infty}\mathcal{I}_{\rm q}/\alpha$. Then, $\kappa_{\infty}\mathcal{I}_{\rm q}/\alpha\times{\rm Reg}$ is not maximized at an $\alpha$ such that $\kappa_{\infty}$ is too close to 1. But we want $\kappa_{\infty}$ to be close enough to 1, hence the regularization term ${\rm Reg}$ should satisfy ${\rm Reg}\simeq 0$ only in the narrow region of $\alpha$ satisfying $\kappa_{\infty}\simeq 1$. We employ the following regularization function satisfying those requirements:
@@
Actually, when we employ the parameter $\delta$ closer to 1, the regularization function becomes sharper around $\alpha$ such that $\kappa_{\infty}\simeq 1$. Hence, we arrive at the objective function given in Eq. (35). In the numerical simulation shown later, we choose $\delta=0.95$.
@@ 3.5 Statistical properties of the estimator
where normal-$\to$ means the convergence in probability as $N\to\infty$. In addition, the random variable $\alpha_{k}(\bm{X}_{k-1})$ converges to a constant $\alpha_{k}^{*}$ in the sense of probability as follows
@@
where $\mathcal{N}(0,1-\langle\mathcal{O}\rangle^{2})$ denotes a centered normal distribution with variance $1-\langle\mathcal{O}\rangle^{2}$, and normal-$\to$ means the convergence in distribution as $N$ increases.
@@
Importantly, the most right hand side in Eq. (45) matches the upper bound in Theorem 1, where we use $N^{*}_{\rm q}~{}(\geq\alpha_{\rm B})$ queries to $A$ and $A^{\dagger}$ in total. In the next section, we demonstrate that the approximation errors are sufficiently small for almost all $\theta^{*}$, via an appropriate choice of $\{D_{k}\}_{k=1}^{M-1}$; that is, together with the asymptotic normality, our estimator nearly achieves the ultimate precision given by the inverse of total quantum Fisher information (22).
@@ 4 Numerical Experiment
4.1 Asymptotic properties with respect to the number of measurement
@@
Figure 2: The estimation error of $\braket{\mathcal{O}}=\cos{\theta^{*}}$ and the total number of queries for several target values (a)–(f). The root mean squared error (RMSE) of the estimator $\cos{\hat{\theta}}$ is evaluated using 300 trials. Since $N_{\rm q}$ in Eq. (49) is a random variable in our method, the RMSE is plotted as a function of the maximum value of $N_{\rm q}$ in 300 trials. The blue solid and green dashed lines represent the asymptotic values of CCR/QCR bounds obtained from the corresponding classical/quantum Fisher information $\mathcal{I}^{*}_{\rm c/q,tot}(\theta^{*})$, respectively. The orange dash-dotted and the gray dotted lines represent the Heisenberg-limited scaling and the precision limit derived in Theorem 1, respectively.
@@
Figure 3: The probability density of the maximum likelihood estimator $\cos{\hat{\theta}_{M}}$ in the case of $M=8$. The blue and orange plots correspond to the results with $N=50$ and $N=500$, respectively. The histograms represent the empirical probability density over 3000 trials, and the solid lines denote the probability density of the centered normal distribution with the variance corresponding to the asymptotic CCR bound. The upper right panels show the same plots with much wider range in the horizontal axis to show the existence of outliers. The larger $N$ becomes, the sharper the density gets around the target value, which demonstrates the consistency in Theorem 3. Also, the outliers vanish in large $N$; thus the empirical density is nearly identical to the normal distribution, which is exactly the asymptotic normality in Theorem 4.
@@
where ##\## is the total number of trials, and $\hat{\theta}[i]$ denotes the estimate of $i$th trial. In the following, we use #$\#=300$ samples to evaluate the RMSE. We assume the depolarization noise with the parameter $p_{\rm q}=0.995$. The number of qubits is set to 20, which corresponds to the system dimension $d=2^{20}$. We also set the number of measurements as $N=500$ for each circuit. To obtain the maximum likelihood estimates, we used a modified brute force method, in which the search domain becomes narrowed as the measurement process proceeds. The amplified level $\alpha_{k}$ of the $k$th measurement process $(k=2,...,M)$ is determined by solving the optimization problem (35), where the maximum likelihood estimate $\hat{\theta}_{k-1}$ has been obtained at the $(k-1)$th step. Also, in this work, the optimization range is chosen as
@@
Figure 2 shows the relationship between the RMSE Eq. (47) and the total number of queries $N_{\rm q}$ calculated as
@@
The red dots indicate the RMSE with $M=3,4,\cdots,12$ from left to right in each panel. Since $N_{\rm q}$ is a random variable due to the randomness of $\alpha_{k}$, we plot the RMSE as a function of the maximum value of $N_{\rm q}$ in #$\#=300$ trials, and thus the RMSE dots are overestimated with respect to the number of queries. In addition, the classical Cramér-Rao (CCR) lower bound and the quantum Cramér-Rao (QCR) lower bound are depicted with the solid blue and dashed green lines, respectively; these are calculated by substituting the true value $\theta^{*}$ and the asymptotic sequence $\{\alpha_{k}^{*}\}_{k=1}^{M}$ into $\mathcal{I}^{*}_{\rm c/q,tot}\left(\theta^{*}\right)$ in Eqs. (42) and (43). Note that the difference between the bounds obtained from the #$\#=300$ average of the total Fisher information (not shown in the figure), $\mathcal{I}_{\rm c/q,tot}\left(\theta^{*}\right)$, and the asymptotic CCR/QCR bounds from $\mathcal{I}^{*}_{\rm c/q,tot}\left(\theta^{*}\right)$ can be negligible, which clearly supports Eq. (41). The orange dash-dotted and gray dotted lines represent the Heisenberg-limited (HL) scaling ${\rm RMSE}=O(1/N_{\rm q})$ and the precision limit given by Eq. (22) with $1-\langle\mathcal{O}\rangle^{2}$, respectively.
Several important features are observed. First, the CCR and QCR bounds are close to each other; the first part of Eq. (45) holds. Note that this good approximation holds even when the target value $\theta^{*}/2\pi$ is a rational number as shown in the panel (d); actually, we will see in the next subsection that this approximation (i.e., the first part of (45)) holds for almost all $\theta^{*}$ in the numerical simulation. We then find that the RMSE almost achieves the CCR and accordingly QCR bounds in all the cases of six target values. That is, $N=500$ is sufficient to obtain the asymptotic normality (4) in the chosen noise condition. Moreover, the empirical probability density of our estimator shown in Fig. 3 converges to the normal distribution whose variance corresponds to the asymptotic CCR bound as $N$ increases, which directly demonstrates the asymptotic normality (4). Recall that this key property holds because the amplified levels are almost identical to $\{\alpha_{k}^{*}\}_{k=1}^{M}$ due to Theorem 3, which is demonstrated in Fig. 7 in Appendix E.

### v4

@@ 3.4 The proposed adaptive estimation algorithm
Our algorithm is based on the maximum likelihood estimation method for the random variables subjected to the probability distribution (3.3). An overview is shown in Fig. 1. The algorithm is composed of the following five procedures.
@@
    Set the first amplified level $\alpha_{1}=1$, and set the number of measurements $N$ and the number of steps $M$ as some natural numbers.
@@
    If the amplified level $\alpha_{k}$ is even, then we measure $N$ copies of the quantum state $\rho_{\mathcal{D}}(\alpha_{k};\theta^{*})$ of Eq. (15) with the POVM (3.3). If $\alpha_{k}$ is odd, we perform the POVM (3.3) to each of $N$ copies of $\rho_{\mathcal{D}}(\alpha_{k};\theta^{*})$. We write the total number of hitting "1" as $x^{(k)}\in\{0,1,\cdots,N\}$. Here, $x^{(k)}$ is a realization of the Binomial random variable
@@
    Determine $\alpha_{k+1}$ for the next measurement based on the current estimate $\hat{\theta}_{k}$ by numerically solving the following optimization problem:
@@
    where $D_{k}\subset\mathbb{N}$ is a (finite) optimization range and $\delta\in(0,1]$ is a parameter for regularization.
  • (v)
    Repeating the above procedure (ii)–(iv) for $k=1,2,...,M$, we obtain a final estimate $\hat{\theta}_{M}$. (We omit the optimization in (iv) when $k=M$.)
We make some remarks on the proposed algorithm. First of all, we assume that the probability of depolarization $1-p_{\rm q}$ is identified by some experiments on the quantum device to be used in advance and the calibration on $p_{\rm q}$ is perfect. The impact of calibration errors of $p_{\rm q}$ on this algorithm is analyzed numerically in Appendix F. Unlike Bayesian estimation, our algorithm does not require any prior information on the target value.
Then, we note that the measurement $\bm{x}_{k}$ in (iii) is the collection of all the results via the measurements with $\alpha_{l}~{}\forall l=1,2,\cdots,k$. As well as the method in Section 2, the ambiguity in estimating $\theta^{*}$, which arises from the periodic nature of $\mathbf{P}_{\mathcal{D}}(\alpha_{k};\theta^{*})$, can be eliminated by the product of likelihoods with different resolutions. Because the amplified level $\alpha_{k}$ depends on the previous results $\bm{x}_{k-1}$ (except for the case $k=1$), the likelihood function $\mathcal{L}_{k}(\theta;\bm{x}_{k})$ has a hierarchical structure, unlike the case of Eq. (6).
The optimization problem (35) comes from the third requirement described in the beginning of this subsection; we will provide a detailed explanation in the next two paragraphs. Since the elements of $D_{k}$ are integers, the optimization problem can be efficiently solved using a simple brute force method. Also, it is obvious that the optimization does not yield an amplified level such that $\sin{(\alpha\hat{\theta}_{k})}\simeq 0$, and thus our method can avoid vanishing classical Fisher information regardless of the target value. It should be noted that this algorithm finally outputs an estimate $\hat{\theta}_{M}$ together with the realized amplified levels $\{\alpha_{k}\}_{k=1}^{M}$ in a single run which uses $N_{\rm q}=N\sum_{k=1}^{M}\alpha_{k}$ queries to $A$ and $A^{\dagger}$ in total.
@@
The optimization presented above has the following issue that forced us to introduce a regularization function. That is, if $\kappa_{\infty}$ is too close to one, the probability $1-\mathbf{P}_{\mathcal{D}}(\alpha;\theta^{*})$ becomes small for such an $\alpha$ when $d\gg 1$, meaning that we need a large number of shots $N$ to have a good estimate. To circumvent this issue, we introduce a regularization function ${\rm Reg}(\alpha;\hat{\theta}_{k})$ such that ${\rm Reg}\simeq 0$ at $\kappa_{\infty}\simeq 1$ and take the objective function $\kappa_{\infty}\mathcal{I}_{\rm q}/\alpha\times{\rm Reg}$ rather than $\kappa_{\infty}\mathcal{I}_{\rm q}/\alpha$. Then, $\kappa_{\infty}\mathcal{I}_{\rm q}/\alpha\times{\rm Reg}$ is not maximized at an $\alpha$ such that $\kappa_{\infty}$ is too close to 1. But we want $\kappa_{\infty}$ to be close enough to 1, hence the regularization term ${\rm Reg}$ should satisfy ${\rm Reg}\simeq 0$ only in the narrow region of $\alpha$ satisfying $\kappa_{\infty}\simeq 1$. We employ the following regularization function satisfying those requirements:
@@
When we employ the parameter $\delta$ closer to 1, the regularization function becomes sharper around $\alpha$ such that $\kappa_{\infty}\simeq 1$. Hence, we arrive at the objective function given in Eq. (35). In the numerical simulation shown later, we choose $\delta=0.95$.
@@ 3.5 Statistical properties of the estimator
where $\to$ means the convergence in probability as $N\to\infty$. In addition, the random variable $\alpha_{k}(\bm{X}_{k-1})$ converges to a constant $\alpha_{k}^{*}$ in the sense of probability as follows
@@
where $\mathcal{N}(0,1-\langle\mathcal{O}\rangle^{2})$ denotes a centered normal distribution with variance $1-\langle\mathcal{O}\rangle^{2}$, and $\to$ means the convergence in distribution as $N$ increases.
@@
Note that this can be rewritten as $N^{*}_{\rm q}=\sum_{l=1}^{M^{\prime}}\alpha^{\prime}_{l}$ for $M^{\prime}=NM$ and $\alpha^{\prime}_{l}=\alpha^{*}_{k}$ if $(k-1)N+1\leq l\leq kN$ in the notation of Theorem 1. Importantly, the most right hand side in Eq. (45) matches the upper bound in Theorem 1, where we use $N^{*}_{\rm q}~{}(\geq\alpha_{\rm B})$ queries to $A$ or $A^{\dagger}$ in total. In the next section, we demonstrate that the approximation errors are sufficiently small for almost all $\theta^{*}$, via an appropriate choice of $\{D_{k}\}_{k=1}^{M-1}$; that is, together with the asymptotic normality, our estimator nearly achieves the ultimate precision given by the inverse of total quantum Fisher information (22).
3.6 Summary of the proposed protocol
Here we summarize the notable feature of the proposed adaptive method, with remarks on which part is theoretically guaranteed and which part is left for numerical verification.
The goal is to ultimately estimate $\langle\mathcal{O}\rangle=\cos\theta^{*}$, for the given $n$-qubit implementable operators $A$ and $\mathcal{O}$; note that, likewise the Grover operator, the implementability of these operators does not mean that all the operator components are known and accordingly $\langle\mathcal{O}\rangle$ is directly computable. The noise parameter $p_{\rm q}$ are assumed to be known. Another important assumption is that $d=2^{n}\gg 1$. We do not need other prerequisite for the target system.
The estimation procedure is illustrated in Fig. 1, where the amplification level $\alpha_{k}$, which determines the number of querying $Q$ and the type of POVM, is updated via Eq. (35) at each step $k$ (each step contains $N$ shots). This optimization (35) is the maximization of classical Fisher information per queries, $\mathcal{I}_{\rm c}/\alpha$, reflecting the goal to achieve quantum Fisher information $\mathcal{I}_{\rm q}$ at each step and $\alpha_{k}\to\alpha_{\rm B}:={\rm argmax}_{\alpha\in\mathbb{N}}~{}\mathcal{I}_{\rm q}(\alpha)/\alpha$. Here, $\alpha_{\rm B}$ provides the best coherent use of noisy $A$ or $A^{\dagger}$ in estimating $\theta^{*}$, which is proved in Theorem 1. From the analysis of our POVM (3.3) in Theorem 2, the objective function $\mathcal{I}_{\rm c}/\alpha$ can be written as $\kappa_{\infty}\mathcal{I}_{\rm q}/\alpha$, where $\kappa_{\infty}$ is a bounded periodic function regarding $\alpha$. Then, to achieve the above goal, the following properties should be satisfied: (i) max $\kappa_{\infty}\mathcal{I}_{\rm q}/\alpha$ in a given optimization range $D_{k}~{}(\ni\alpha)$ is close to max $\mathcal{I}_{\rm q}/\alpha$ in $D_{k}$ and (ii) $\kappa_{\infty}\approx 1$ (i.e., $\mathcal{I}_{\rm c}\approx\mathcal{I}_{\rm q}$) at the maximum point. Note that we gradually expand the range of $D_{k}$ as $k$ increases to eliminate the estimation ambiguity; see Section 2. In Section 3.4, we provide an intuitive explanation (not a rigorous proof) that the properties (i) and (ii) hold, and furthermore, these properties will be numerically verified in the next section.
On the other hand, Theorems 3 and 4 guarantee the statistical properties of the estimator. More precisely, Theorem 3 guarantees the consistency, stating that the estimate approaches the true value with high probability as the number of shots $N$ increases. Also, Theorem 4 guarantees the asymptotic normality, stating that the estimator asymptotically achieves the classical Cramér-Rao lower bound $1/\mathcal{I}^{*}_{\rm c,tot}(\theta^{*})$. Here, the (asymptotic) total classical/quantum Fisher information $\mathcal{I}^{*}_{\rm c,tot}(\theta^{*})$ and $\mathcal{I}^{*}_{\rm q,tot}(\theta^{*})$ are defined in Eqs. (42) and (43), respectively. Therefore, if the nearly optimal sequence $\alpha_{k}$ is properly chosen, the estimator achieves the estimation precision $1/\mathcal{I}^{*}_{\rm c,tot}(\theta^{*})\approx 1/\mathcal{I}^{*}_{\rm q,tot}(\theta^{*})$, and furthermore, it also achieves the precision ${\alpha_{\rm B}}/{(N^{*}_{\rm q}\mathcal{I}_{\rm q}(\alpha_{\rm B}))}$ for the (asymptotic) total queries $N^{*}_{\rm q}$ in Eq. (46), when the iteration $k$ is sufficiently large. This matches the precision limit in Theorem 1 (more precisely, Eq. (22)). We numerically find that this happens in a moderate number of shots $N$, as will be demonstrated in Figs. 2, 3 and 5.
@@ 4 Numerical Experiment
4.1 Asymptotic properties with respect to the number of measurements
@@
Figure 2: The estimation error of $\braket{\mathcal{O}}=\cos{\theta^{*}}$ and the total number of queries for several target values (a)–(f). The x-axis shows the total number of queries to the state preparation $A$ or $A^{\dagger}$ in a single trial of our method. The root mean squared error (RMSE) of the estimator $\cos{\hat{\theta}}$ is estimated by the averaged value over 300 trials, as in Eq. (4.1). Since the number of total queries in a single trial of our method varies stochastically, the corresponding x-axis value of the RMSE dot is taken as the maximum value of total queries in the 300 trials. The blue solid and green dashed lines represent the asymptotic values of CCR/QCR bounds obtained from the corresponding classical/quantum Fisher information $\mathcal{I}^{*}_{\rm c/q,tot}(\theta^{*})$, respectively. The orange dash-dotted and the gray dotted lines represent the Heisenberg-limited scaling and the precision limit derived in Theorem 1 (in the case of $N_{\rm q}>\alpha_{\rm B}=199$), respectively.
@@
Figure 3: The probability density of the maximum likelihood estimator $\cos{\hat{\theta}_{M}}$ in the case of $M=8$. The blue and orange plots correspond to the results with $N=50$ and $N=500$, respectively. The histograms represent the empirical probability density over 3000 trials, and the solid lines denote the probability density of the centered normal distribution with the variance corresponding to the asymptotic CCR bound. The upper right panels show the same plots (but the x-axis represents the absolute error) with much wider range in the horizontal axis to show the existence of outliers. The larger $N$ becomes, the sharper the density gets around the target value, which demonstrates the consistency in Theorem 3. Also, the outliers vanish in large $N$; thus the empirical density is nearly identical to the normal distribution, which is exactly the asymptotic normality in Theorem 4.
@@
where ##\## is the total number of trials, and $\hat{\theta}[i]$ denotes the estimate of $i$th trial. In the following, we use #$\#=300$ samples to evaluate the RMSE. We assume the depolarization noise with the parameter $p_{\rm q}=0.995$. The number of qubits is set to 20, which corresponds to the system dimension $d=2^{20}$. We also fix the number of measurements as $N=500$ for each circuit. To obtain the maximum likelihood estimates, we used a modified brute force method, in which the search domain becomes narrowed as the measurement process proceeds. The amplified level $\alpha_{k}$ of the $k$th measurement process $(k=2,...,M)$ is determined by solving the optimization problem (35), where the maximum likelihood estimate $\hat{\theta}_{k-1}$ has been obtained at the $(k-1)$th step. Also, in this work, the optimization range is chosen as
@@
Figure 2 shows the relationship between the RMSE Eq. (4.1) and the total number of queries $N_{\rm q}$ calculated as
@@
Here, we remark that $N_{\rm q}$ can be rewritten as $N_{\rm q}=\sum_{l=1}^{M^{\prime}}\alpha_{l}^{\prime}$ for $M^{\prime}=NM$ and $\alpha^{\prime}_{l}=\alpha_{k}$ if $(k-1)N+1\leq l\leq kN$ in the notation of Theorem 1. The red dots indicate the RMSE with $M=3,4,\cdots,12$ from left to right in each panel. Since $N_{\rm q}$ is a random variable due to the randomness of $\alpha_{k}$, we plot the RMSE as a function of the maximum value of $N_{\rm q}$ in #$\#=300$ trials, and thus the RMSE dots are overestimated with respect to the number of queries. In addition, the classical Cramér-Rao (CCR) lower bound and the quantum Cramér-Rao (QCR) lower bound are depicted with the solid blue and dashed green lines, respectively; these are calculated by substituting the true value $\theta^{*}$ and the asymptotic sequence $\{\alpha_{k}^{*}\}_{k=1}^{M}$ into $\mathcal{I}^{*}_{\rm c/q,tot}\left(\theta^{*}\right)$ in Eqs. (42) and (43). Note that the difference between the bounds obtained from the #$\#=300$ average of the total Fisher information (not shown in the figure), $\mathcal{I}_{\rm c/q,tot}\left(\theta^{*}\right)$, and the asymptotic CCR/QCR bounds from $\mathcal{I}^{*}_{\rm c/q,tot}\left(\theta^{*}\right)$ can be negligible, which clearly supports Eq. (41). The orange dash-dotted and gray dotted lines represent the Heisenberg-limited (HL) scaling ${\rm RMSE}=O(1/N_{\rm q})$ and the precision limit given by Eq. (22) with $1-\langle\mathcal{O}\rangle^{2}$, respectively.
Several important features are observed. First, the CCR and QCR bounds are close to each other; the first part of Eq. (45) holds. Note that this good approximation holds even when the target value $\theta^{*}/2\pi$ is a rational number as shown in the panel (d); actually, we will see in the next subsection that this approximation (i.e., the first part of (45)) holds for almost all $\theta^{*}$ in the numerical simulation. We then find that the RMSE almost achieves the CCR and accordingly QCR bounds in all the cases of six target values. That is, $N=500$ is sufficient to obtain the asymptotic normality (4) in the chosen noise condition. We here remark that by taking the number of shots $N$ (originally, we fix $N=500$ for all iteration steps) as a linear function $N_{k}$ of the iteration step $k$, as was done in Ref. [46, 63], the RMSE of our method can nearly achieve the QCR bound even for a regime with smaller total queries ($N_{\rm q}\sim O(10^{2})$). Moreover, the empirical probability density of our estimator shown in Fig. 3 converges to the normal distribution whose variance corresponds to the asymptotic CCR bound as $N$ increases, which directly demonstrates the asymptotic normality (4). Recall that this key property holds because the amplified levels are almost identical to $\{\alpha_{k}^{*}\}_{k=1}^{M}$ due to Theorem 3, which is demonstrated in Fig. 7 in Appendix E.